OpenStudy (anonymous):

a carpenter is building a rectangular room with a fixed perimeter of 92ft. what demensions would yeild the maximum area? i got 529ft is that correct?

OpenStudy (anonymous):

92=2l+2w 46=l+w 23^2=529 I dont know if I done this right or not but thats the way i was working it

OpenStudy (mertsj):

A square will give the largest area. If the perimeter is 92 each side is 23 so the area is 23^2

OpenStudy (anonymous):

Let the width be w Let the length be l 2(w+l) = 92 w = 46 - l l(46 - l) 23^2 = 529

OpenStudy (mertsj):

The correct way to do it is to say the sides of the rectangle are x and 46-x so the area is x(46-x) Take the first derivative, set it to 0 and solve for x.

OpenStudy (anonymous):

i think i did this wrong, the length that would yield the maximum area is what? would it still be 529? or do i need to start over

OpenStudy (mertsj):

It would be 23

OpenStudy (anonymous):

would you mind showing me how you got that

OpenStudy (mertsj):

Are you taking calculus?

OpenStudy (anonymous):

You don't need calculus because the equation is a parabola, the maximum is the vertex

OpenStudy (anonymous):

algebra and my teacher thinks i know how to do all this stuff lol I DONT

OpenStudy (mertsj):

That is correct but it is just as easy to take the derivative if she is taking calculus.

OpenStudy (anonymous):

i wished i didnt have to take any math lol my math teacher in high school made me not like math

OpenStudy (anonymous):

ok so now i got to find the width

OpenStudy (anonymous):

area = x(46-x) = -x^2+46x The x coordinate for the vertex of a parabold is -b/2a which equals -46/-2 = 23

OpenStudy (mertsj):

ok then. You should write that the area is x(46-x) or 46x-x^2. That is a parabola. The x coordinate of its vertex is -b/2a or -46/-2 or 23. So one side is 23 and the other side is 46-23 which is also 23. So it is a square---a 23 by 23 square.

OpenStudy (anonymous):

so the length would be my original answer of 529?

OpenStudy (anonymous):

no its l=23 w=23 maximum =529